Jefferson Lab HALL A Logbook for September 99 - All EXCEPT Automatic Entries

  • Main INDEX

    Thursday

  • 27368 09/30/99 23:45 M. Liang FPP HV off. Coda crash
  • 27366 09/30/99 23:27 M. Liang Started Run for LD2 with radiator in
  • 27365 09/30/99 23:13 Perdrisat target incident (harmless)
  • 27356 09/30/99 22:33 M. Liang MCC testing high current
  • 27354 09/30/99 21:57 M. Liang He flow back to normal
  • 27353 09/30/99 21:36 Churchwell HIRESXT20 problem
  • 27352 09/30/99 20:24 Churchwell Helium flow alarms in spectrometers
  • 27351 09/30/99 19:28 M. Liang He flow of Quads are too low
  • 27350 09/30/99 17:27 R. Michaels super datamon
  • 27348 09/30/99 16:34 krishniw FPP HV change
  • 27347 09/30/99 16:30 krishniw shift activity
  • 27342 09/30/99 09:40 gilman large FPP th/phi front events
  • 27340 09/30/99 07:39 Nilanga Optics checks
  • 27339 09/30/99 02:47 x jiang magnet screen at run 1899
  • Wednesday

  • 27337 09/29/99 23:55 R Roche shift summary 16:00-24:00
  • 27323 09/29/99 18:51 R. Michaels Dead Bell in Datamon
  • 27318 09/29/99 18:26 R Roche update on spectrometers
  • 27306 09/29/99 14:43 gilman/mitch more on FPP hardware
  • 27303 09/29/99 12:02 gilman/diete fpp hardware check
  • 27300 09/29/99 10:03 saha Arc Energy measurement, Sept 22, 1999
  • 27299 09/29/99 07:52 x jiang maintenance day.
  • 27298 09/29/99 00:02 R Roche Target Status 16:00-24:00
  • Tuesday

  • 27296 09/28/99 16:28 W.Xu Target situation in Maintenance day
  • 27292 09/28/99 16:22 bogdanw HA PMT
  • 27291 09/28/99 12:15 deurpam eP offline analysis result
  • 27290 09/28/99 11:12 gilman more background study, with cosmics
  • 27287 09/28/99 00:20 x jiang t 0:15, ended last run, run 1876 of this shift.
  • Monday

  • 27285 09/27/99 23:58 M. Liang Swing shift summary
  • 27283 09/27/99 23:16 M. Liang Beam off
  • 27280 09/27/99 23:06 M. Liang Currnt dropping towards 3muA
  • 27276 09/27/99 18:50 M. Liang radiator out, 2H target in
  • 27270 09/27/99 15:54 Huber Day Shift Summary
  • 27264 09/27/99 13:56 Huber Beam Polarization Changed
  • 27262 09/27/99 13:26 Eugene Moller results
  • 27261 09/27/99 10:44 R. Michaels How to diagnose DAQ (NEW)
  • 27256 09/27/99 07:56 Z. Chai Owl Shift Summary
  • 27253 09/27/99 07:18 Z. Chai Beam Current
  • 27239 09/27/99 01:38 Z. Chai VDC Status
  • 27238 09/27/99 00:06 R. Michaels trigger checks
  • Sunday

  • 27236 09/26/99 23:39 gilman shift summary - 9/26/1999 swing shift
  • 27232 09/26/99 20:23 gilman beam current since noon fix
  • 27231 09/26/99 19:58 R. Michaels RE: DAQ hangup
  • 27229 09/26/99 19:33 gilman coda crash?
  • 27216 09/26/99 16:01 Huber Day shift summary
  • 27207 09/26/99 10:41 Huber Oscillation in Beam Intensity
  • 27201 09/26/99 07:55 Z. Chai Owl Shift Summary
  • Saturday

  • 27185 09/25/99 23:23 C.F. Perdris chopper ok
  • 27184 09/25/99 23:17 C.F. Perdris background continued
  • 27183 09/25/99 22:08 meziani no beam
  • 27182 09/25/99 20:25 meziani Beam tune
  • 27181 09/25/99 20:08 meziani Beam spot size
  • 27180 09/25/99 19:52 meziani Beam back
  • 27178 09/25/99 19:01 Z.-E. Mezian No beam for tuning
  • 27176 09/25/99 17:15 Z.-E. Mezian Recesiation-followup
  • 27174 09/25/99 16:50 Z.-E. Mezian Recesiation of the crystal
  • 27172 09/25/99 15:55 Huber Day shift summary
  • 27168 09/25/99 13:16 gilman statistics goals / run times
  • 27166 09/25/99 10:02 gilman background thoughts
  • 27162 09/25/99 07:56 S. Strauch beam current overview
  • 27161 09/25/99 07:55 Z. Chai Owl Shift Summary
  • 27156 09/25/99 06:35 Z. Chai H-Arm Z Target
  • 27137 09/25/99 00:00 M. Liang Swing shift summary
  • Friday

  • 27134 09/24/99 23:49 R. Michaels RE: CODA problems
  • 27133 09/24/99 23:19 M. Liang Pedestal runs
  • 27130 09/24/99 23:09 M. Liang coda problem
  • 27123 09/24/99 21:47 C.F. Perdris background analysis
  • 27122 09/24/99 20:59 M. Liang beam current drop fast
  • 27121 09/24/99 20:56 M. Liang shift check list
  • 27114 09/24/99 16:55 M. Liang radiator in,2H target in
  • 27110 09/24/99 16:24 M. Liang radiator out
  • 27108 09/24/99 16:01 Huber Day Shift Summary
  • 27104 09/24/99 13:36 R. Michaels ROC14 and other DAQ crashes
  • 27100 09/24/99 11:47 Eugene Moller results
  • 27091 09/24/99 07:54 Z. Chai Owl Shift Summary
  • 27081 09/24/99 04:31 Z. Chai ROC 14 problem
  • 27080 09/24/99 04:03 Z. Chai Spot Check
  • 27078 09/24/99 03:12 Z. Chai Spot Check
  • 27066 09/24/99 00:04 M. Liang swing shift summary
  • Thursday

  • 27064 09/23/99 23:57 punjabi status of detectors when started after energy change
  • 27063 09/23/99 23:46 M. Liang beam away
  • 27061 09/23/99 23:26 M. Liang 2H target in position
  • 27055 09/23/99 22:53 M. Liang 6% radiator in
  • 27054 09/23/99 22:43 deurpam eP online result.
  • 27053 09/23/99 21:21 M. Liang, A. ep measurement started
  • 27052 09/23/99 17:42 M. Liang no beam
  • 27051 09/23/99 16:21 gilman shift summary
  • 27049 09/23/99 15:45 M Jones Wien filter
  • 27048 09/23/99 14:36 Jack Hadron HVAC is running.
  • 27047 09/23/99 09:09 deurpam eP measurement result
  • 27038 09/23/99 07:59 S. Strauch Owl shift summary
  • 27025 09/23/99 05:25 S. Strauch low beam current
  • 26996 09/23/99 00:02 M. Liang Swing shift activities
  • Wednesday

  • 26993 09/22/99 23:55 M. Liang radiator out
  • 26989 09/22/99 23:22 M. Liang 6% radiator in
  • 26988 09/22/99 23:20 M. Liang 1.5"+3" Carbon door in
  • 26958 09/22/99 17:42 M. Liang radiator out
  • 26957 09/22/99 17:31 M. Liang Earm set. start arc measurement
  • 26956 09/22/99 17:15 M. Liang Controlled access finished.
  • 26955 09/22/99 17:12 M. Liang 1.5" carbon door in. empty target in position.
  • 26954 09/22/99 16:52 M. Liang acitivities: finished C(e,e), moving E,Harm
  • 26947 09/22/99 16:15 gilman 9-22 day shift summary
  • 26942 09/22/99 13:54 deurpam eP energy measurement
  • 26941 09/22/99 12:20 gilman e-arm angle change not documented last evening
  • 26933 09/22/99 09:14 M JOnes Checks of harps
  • 26930 09/22/99 08:35 M Jones Checked fpp current and gas flow
  • 26929 09/22/99 08:02 S. Strauch owl shift summary
  • 26910 09/22/99 00:07 M. Liang swing shift summary
  • Tuesday

  • 26909 09/21/99 23:55 M. Liang ROC 14 needs reboot
  • 26907 09/21/99 23:18 M. Liang Beam away for 20 minutes
  • 26902 09/21/99 21:29 V. Punjabi difference in number of zeros between LH2 and LD2 runs
  • 26899 09/21/99 20:09 M. Liang low efficiency due to background?
  • 26896 09/21/99 20:01 M. Liang gammap_online.kumac
  • 26892 09/21/99 18:22 M. Liang low effieciency
  • 26889 09/21/99 17:49 M. Liang beam back
  • 26888 09/21/99 17:44 M. Liang low effieciency in FPP and VDC?
  • 26884 09/21/99 16:04 gilman,chen, day shift summary
  • 26875 09/21/99 13:32 gilman polarized beam current
  • 26868 09/21/99 11:44 gilman hadron spectrometer adjusted
  • 26867 09/21/99 11:31 Eugene Moller measurement
  • 26866 09/21/99 11:28 jpchen cryotarget ioc reboots
  • 26865 09/21/99 11:25 jpchen LD2 is set to be ready for beam
  • 26856 09/21/99 08:13 Fox,Jones,Ga Shift Summary - 09/21/99 - Owl
  • 26850 09/21/99 07:22 Fox FPP voltages raised again
  • 26849 09/21/99 07:21 Fox Configuration for Q2=0.3486 (e,ep) elastic data
  • 26844 09/21/99 06:03 Fox Beam Spot off center for runs 1700 thru 1702, fixed during 1702
  • 26840 09/21/99 05:18 Fox Errata for Start run entry for run 16999
  • 26835 09/21/99 04:12 M Jones Downloaded trigger for 0.514 proton momentum
  • 26834 09/21/99 04:12 Fox Settings for elastic (e,p) on H, Q2=0.246887
  • 26833 09/21/99 04:09 Fox Raised FPP HV due to lowish efficiency
  • 26825 09/21/99 02:07 Fox Beam spot is (2,4) instead of (2,2)
  • 26817 09/21/99 00:58 M Jones checked gas flow
  • 26816 09/21/99 00:57 M Jones checked fpp currents
  • Monday

  • 26814 09/20/99 23:56 M. Liang Carbon door closed, ready to take ep elastic data
  • 26813 09/20/99 21:27 M. Liang optimistc estimate: beam by midnight
  • 26812 09/20/99 19:06 M. Liang a piece of paper found in Hall B line
  • 26811 09/20/99 18:18 M. Liang No beam for at least next 3 hours. MCC need access to the tunnel
  • 26810 09/20/99 16:52 M. Liang beam permit
  • 26809 09/20/99 16:48 R Roche, M. No beam
  • 26807 09/20/99 15:17 krishniw Receciation
  • 26806 09/20/99 14:53 gilman beam imminent, carbon doors reset
  • 26803 09/20/99 09:19 gilman gas for phototubes
  • Sunday

  • 26798 09/19/99 22:01 gilman gas to phototubes
  • 26795 09/19/99 16:09 kuss target chamber vacuum
  • 26791 09/19/99 15:11 Calarco shift summary and prognosis
  • 26790 09/19/99 12:35 kuss still no beam
  • Saturday

  • 26787 09/18/99 20:40 Perdrisat update on beam
  • 26786 09/18/99 19:53 Perdrisat beam expectation
  • 26785 09/18/99 18:32 Perdrisat carbon door
  • 26782 09/18/99 16:49 Calarco shift summary
  • 26780 09/18/99 10:57 M Jones fixed fpp_det.kumac
  • 26779 09/18/99 08:01 Z.-E. Mezian Target chamber vacuum
  • 26777 09/18/99 07:31 R. Michaels detector checks, more cosmics
  • 26774 09/18/99 03:38 R. Michaels status, detector checks
  • 26773 09/18/99 01:02 fox Problem with hadron aerogel ADC 12
  • Friday

  • 26772 09/17/99 23:24 D. Margaziot No beam
  • 26769 09/17/99 22:44 gilman e-arm HV off
  • 26768 09/17/99 16:56 x. jiang checked the radiator movement
  • 26767 09/17/99 16:54 x. jiang fpp and vdc voltages checked.
  • 26766 09/17/99 16:51 x jiang fpp hv channels.
  • Wednesday

  • 26761 09/15/99 12:42 R. Michaels H+,H- rates ok
  • 26760 09/15/99 11:27 R. Michaels end-run H+/- rates strange
  • 26756 09/15/99 00:54 R. Michaels VDC efficiency - cosmics
  • Tuesday

  • 26755 09/14/99 22:47 folts magic
  • 26753 09/14/99 15:22 R. Michaels ROC1 power supply bad
  • 26750 09/14/99 10:07 K. Aniol target temperature excursion
  • Monday

    
    

    Sunday

    
    

    Saturday

    
    

    Friday

    
    

    Thursday

  • 26715 09/09/99 12:27 R. Michaels DAQ mods, etc.
  • Friday

  • 26714 09/03/99 20:20 R. Michaels linux versio - halog
  • 26713 09/03/99 20:19 R. Michaels HP-UX test
  • 26712 09/03/99 20:18 R. Michaels halogentry test
  • 26698 09/03/99 16:11 R. Michaels run 1485 recovered
  • Wednesday

  • 26693 09/01/99 01:09 kuss ARC PMT's
  • 26692 09/01/99 00:03 R Roche Shift summary